A motor boat going downstream overcomes a float at a point A. 60 minutes later it turns and after some time passes the float at a distance of 12 Km from the point A. The velocity of the stream is (assuming constant velocity for the boat in still water).

Grade:11

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Let speed of stream be v and speed of motorboat be u.
Now distance moved by boat in 1 hour=(u+v)*1=u+v
total time is=6/v
speed of boat when it turns back=u-v
=>u+v-6=(u-v)(6/v -1)
solving for v we get, v=3 km/hr [Ans]
Regards,
Thank You,
Soumya Ranjan Mohanty.